OpenStudy (anonymous):

A change jar holds 36 nickels, 30 dimes, and 15 quarters. What is the probability that a nickel will be drawn first and then a dime (without replacement)?

OpenStudy (anonymous):

total number = 45+36 = 81

OpenStudy (anonymous):

P(N, D) = (36/81) x (30/ 80)

OpenStudy (anonymous):

assuming the coins are the same size and all the other things, even though thats not possible

OpenStudy (anonymous):

You have to assume they are blindfolded and wearing thick gloves and cannot gauge the weights of the coins

OpenStudy (anonymous):

thank you for the help :)

OpenStudy (anonymous):

= 1/6